Average Rate of Return Method, Net Present Value Method, and Analysis for a service company

The capital investment committee of Arches Landscaping Company is considering two capital investments. The estimated income from operations and net cash flows from each investment are as follows:

Front-End Loader Greenhouse
Year Income from
Operations
Net Cash
Flow
Income from
Operations
Net Cash
Flow
1 $34,200 $104,000 $72,000 $166,000
2 34,200 104,000 55,000 140,000
3 34,200 104,000 27,000 99,000
4 34,200 104,000 12,000 68,000
5 34,200 104,000 5,000 47,000
Total $171,000 $520,000 $171,000 $520,000

Each project requires an investment of $380,000. Straight-line depreciation will be used, and no residual value is expected. The committee has selected a rate of 10% for purposes of the net present value analysis.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1a. Compute the average rate of return for each investment. If required, round your answer to one decimal place.

Average Rate of Return
Front-End Loader %
Greenhouse %

1b. Compute the net present value for each investment. Use the present value of $1 table above. If required, use the minus sign to indicate a negative net present value.

Front-End Loader Greenhouse
Present value of net cash flow $ $
Amount to be invested $ $
Net present value $ $

2. Prepare a brief report for the capital investment committee, advising it on the relative merits of the two investments.

The front-end loader has a   net present value because cash flows occur   in time compared to the greenhouse. Thus, if only one of the two projects can be accepted, the   would be the more attractive.
Answer #1

Average Rate of Return = Average Income/Average Investment

Average Investment = (380,000+0)/2 = 190,000

Front-end loader= 34,200/190,000

= 18%

Greenhouse = 34,200/190,000

= 18%

Front end loader:

Present value of inflows = 104,000*PVAF(10%, 5 years)

= 104,000*3.79 = $394,160

Less: Amount to be invested = 380,000

NPV = $14,160

Greenhouse:

Present value of inflows = 166,000*0.909 + 140,000*0.826 + 99,000*0.751 + 68,000*0.683 + 47,000*0.621

= $416,514

Less: Amount to be invested = 380,000

NPV = $36,514

Lower NPV

Later

Greenhouse would be more attractive